OpenStudy (anonymous):

Algebra Question what does x=? y = x _________ x-5

OpenStudy (anonymous):

y = -5? i'm guessing. Since you cross out the x in the denominator and numerator.

OpenStudy (anonymous):

this is wrong !!! -5y x = ------ 1 + y

myininaya (myininaya):

i cant read this is it y/x=x/(-5) or y/(x-5)=x/(x-5) ?